thirty kg of nickel that was 38% pure was mixed with a nickel alloy that was 28%pure. how many kg of each kind were used to produce and alloy that was 32% pure.
Answer by checkley77(12844)

.38*30+.28x=.32(30+x)
11.4+.28x=9.6+.32x
.28x-.32x=9.6-11.4
-.04x=-1.8
x=-1.8/-.04
x=45 kg. of 28% alloy ans.
Proof:
.38*30+.28*45=.32(30+45)
11.4+12.6=.32*75
24=24